OpenStudy (anonymous):

The ocean club collects 24.93 kg of trash on the first weekend of the summer. The club collects 18.3 kg the next weekend. How much more trash is collected the first weekend? Give your answer with the appropriate precision. A. 6.7 kg B. 6.65 kg C. 6.63 kg D. 6.6 kg

OpenStudy (anonymous):

Let x more thrash collected so, 18.3+x=24.93 Find x from here.

OpenStudy (destinymasha):

subtract 18.3 from 24.93

OpenStudy (anonymous):

I taught him in PMs, anymore problems?

OpenStudy (destinymasha):

= 6.63

OpenStudy (destinymasha):

C

OpenStudy (anonymous):

^ Thanks

OpenStudy (anonymous):

Michael bikes 36 miles last week and 38.4 miles this week. What is the average distance he bikes? Give your answer with the appropriate precision. A. 36 mi B. 36.2 mi C. 37 mi D. 37.2 mi

OpenStudy (anonymous):

Average means just take the sum and divide it by 2! So, whats (36+38.4)/2= ... ?

OpenStudy (destinymasha):

average = add them all up , and divide it by the 3 of numbers there are

OpenStudy (destinymasha):

36 +38.4 = 74.4 divided by 2 = 37.2 [D]

OpenStudy (anonymous):

If I were you Destiny, I would let him calculate it instead of just giving the answer (just an opinion)

OpenStudy (destinymasha):

sure ......

OpenStudy (anonymous):

Should an estimate or the exact answer be calculated? Then do the appropriate calculation. The sewing club needed 90 yards of lace to finish their various projects. They had 2 rolls of lace that had yards on each roll and 3 rolls with yards of lace on each roll. Did they have enough lace to finish their projects? A. An estimate can be used to know if they have enough lace. 12 • 2 + 24 • 3 = 96 They have at least 96 yards of lace. B. The exact answer is needed to know if they have enough lace. They have yards of lace.

OpenStudy (anonymous):

You need to know the exact number of lace. Estimating would be impossible seeing as there are no yards on each roll given and you would be just guessing at numbers and would not know if they had enough to complete the project.

OpenStudy (anonymous):

get is in b and net is next to the two and bet is by the three

OpenStudy (anonymous):